When a 7.00 g sample of KCN is dissolved in water in a calorimeter that has a total heat capacity of 3.40 kJ⋅K−1, the temperature decreases by 0.370 K. Calculate the molar heat of solution of KCN.
1
Expert's answer
2020-05-19T08:42:32-0400

Moles of KCN=7.00/65=0.107 mol

Total heat capacity=n*C*∆T

3.40=0.107*C*0.370

C=3.40/0.107*0.370

C=85.88 Kj/mol